OpenStudy (anonymous):

Write the equation of the line in point-slope form that passes through the points (-5, -2) and (-4, -5).

OpenStudy (marco26):

Use two-point formula and rearrange the equation in the form y=mx+b

OpenStudy (marco26):

\[\frac{y-y1}{y2-y1}=\frac{x-x1}{x2-x1}\]

OpenStudy (anonymous):

which ones the y's an which ones the x's

OpenStudy (pfenn1):

(x,y)

OpenStudy (pfenn1):

x1 = -5, x2 = -2, y1 = -2, y2 = -5

OpenStudy (anonymous):

im confused

OpenStudy (pfenn1):

plug the values for x1, x2, y1, y2 into the equation that @marco26 gave you above.

OpenStudy (pfenn1):

(x1, y1) = (-5, -2) (x2, y2) = (-4, -5)